• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 09:26:44 on February23,2002

    Entry number 3701

    keyword=target ioc reboot

    At the beginning of shift I noticed that the target alarm was frozen, temperature of target shown on the monitor was 19.33 degree, the plot on loop2 pannel showed a staight line at 19.01 degree, other plots on loop2 pannel were not updating too.

    Around 8:40, when beam went off for controlled access, target temperature on monitor started to drop. I rebooted target ioc, about 5 minutes later target temperature started to come back slowly.

    After ioc reboot, maximum setting for high power heater was 0.00, i set it slowly back to 400W. ~9:10am, target is stable and beeps nicely when beam goes on/off.